A certain drug is found to be effective 60% of the time. Find the probability of succesful treatment in two out of five cases.

.2410

.0400

.1740

.2304

.0500
Answer #1

Answer: .2304

We are given

n = 5

p = 0.6

q = 1 – 0.6 = 0.4

P(X=x) = nCx*p^x*q^(n – x)

P(X=2) = 5C2*0.6^2*0.4^(5 – 2)

P(X=2) = 10*0.6^2*0.4^3

P(X=2) = 0.2304